OpenStudy (anonymous):

simplify the sqaure root of 108

OpenStudy (anonymous):

108=36(3) SQUARE ROOT 36=6 SQUARE ROOT 108=6 times square root of 3

OpenStudy (anonymous):

\[\sqrt{36 * 3}\] \[\sqrt{6*6*3}\] Then the primes \[\sqrt{2*3*2*3*3}\] But I think in your case for simplicity is \[\sqrt{108} \] =\[\sqrt{36*3}\] \[\sqrt{6} \sqrt{6} \sqrt{3}\] \[6 \sqrt{3}\]